# Inequality Restricted Minimum Density Power Divergence Estimation in Panel Count Data

### v1

@@ Abstract
Analysis of panel count data has garnered a considerable amount of attention in the literature, leading to the development of multiple statistical techniques. In inferential analysis, most of the works focus on leveraging estimating equations-based techniques or conventional maximum likelihood estimation. However, the robustness of these methods is largely questionable. In this paper, we present the robust density power divergence estimation for panel count data arising from nonhomogeneous Poisson processes, correlated through a latent frailty variable. In order to cope with real-world incidents, it is often desired to impose certain inequality constraints on the parameter space, giving rise to the restricted minimum density power divergence estimator. The significant contribution of this study lies in deriving its asymptotic properties. The proposed method ensures high efficiency in the model estimation while providing reliable inference despite data contamination. Moreover, the density power divergence measure is governed by a tuning parameter $\gamma$, which controls the trade-off between robustness and efficiency. To effectively determine the optimal value of $\gamma$, this study employs a generalized score-matching technique, marking considerable progress in the data analysis. Simulation studies and real data examples are provided to illustrate the performance of the estimator and to substantiate the theory developed.
@@
Asymptotic property , Density power divergence , Inequality restrictions , Optimal tuning parameter , Robustness
@@
CONTACT Shuvashree Mondal Email: shuvasri29@iitism.ac.in
@@ 1 Introduction
In the study of event history, the types of events that occur repetitively are usually referred to as recurrent events. Some examples of recurrent events refer to hospitalizations of intravenous drug users, multiple instances of specific tumors (Wang et al. 2001), repeated pyogenic infections in patients with inherited disorders (Lin et al. 2000), and claims for vehicle warranties (Kalbfleisch et al. 1991). Panel count data are common when subjects in a study encounter recurrent events observed only at discrete time points instead of being observed continuously in time. They might occur when continuous monitoring of the subjects under study is too expensive or infeasible. As a result, panel count data only provide the number of occurrences of the recurrent events of interest between successive observation times, while the exact occurrence times remain unknown; refer to Huang et al. 2006; Sun and Zhao 2013.
@@
In the literature, it is a common practice to characterize the occurrences of recurrent events by leveraging counting processes. Further, due to the incomplete nature of the panel count data, it is more convenient to model the mean function of the counting processes. Readers may refer to Sun and Kalbfleisch 1995; Wellner and Zhang 2000; Hu et al. 2009 and references therein. In this work, we focus on the analysis of panel count data where each subject under study is exposed to multiple recurrent events characterized by nonhomogeneous Poisson processes. It is a non-stationary counting process with independent increments. In many instances, as it proves challenging to assess all the covariates, a latent frailty variable can be considered that reflects individual-specific heterogeneity typically influencing the frequency of these recurrent incidents; follow Slimacek and Lindqvist 2016. In this work, we develop the inferential analysis of panel count data where each subject under study is exposed to multiple recurrent events characterized by nonhomogeneous Poisson processes correlated through a frailty variable. Based on the counting process, there exists a substantial body of literature concerning inferential studies for panel count data. Amongst them, He et al. 2009 exploited a semi-parametric approach for the analysis of panel count data. Moreover, Sun and Wei 2000 and He et al. 2008 investigated regression analysis based on counting processes concerning univariate and multivariate panel count data, respectively.
@@
Throughout the years, most research on estimation methods for panel count data has been devoted to maximum likelihood estimation (MLE). The major contributions consist of the works by Wellner and Zhang 2007; Zhu et al. 2018; Li et al. 2021; Zeng and Lin 2021. Maximum likelihood estimation is a popular method because of its recognized characteristics, such as asymptotic efficiency, consistency, sufficiency, and invariance. However, with small deviations from the assumed model conditions, those methods often produce biased and inefficient estimates, highlighting the necessity for robust estimation techniques. In this paper, our objective is to perform statistical inference based on multivariate panel count data by applying robust density power divergence estimation methods. The minimum density power divergence estimation (MDPDE) was first proposed by Basu et al. 1998 and subsequently applied in numerous studies. Readers can track the publications of Ghosh and Basu 2013; Basu et al. 2022; Mandal et al. 2023. Recently, Balakrishnan et al. 2019 developed weighted minimum density power divergence estimators to analyze one-shot device testing data, similar to current status data, which is a specific type of panel count data. Upon reviewing the literature, it is found that none of the existing studies on multivariate panel count data generated from the counting process address robust methods for parameter estimation. Hence, we suggest this novel approach to serve the need for a robust estimation technique.
@@
Further, in real-life scenarios, it is often observed that one recurring event takes place more frequently than the other event. Such instances can be accurately characterized by imposing restrictions on the parameter space. Moreover, research on restricted minimum density power divergence estimators has been sparse, with all current studies concentrating solely on equality constraints. For example, Basu et al. 2018; Felipe et al. 2023; Baghel and Mondal 2024 suggested equality constraints in the parameter space of the model parameters and derived the restricted minimum density power divergence estimators. In this study, we broaden the scope of restricted minimum density power divergence estimators by introducing inequality constraints into the parameter space. The principal novelty in this paper lies in deriving the asymptotic distribution of the minimum density power divergence estimators while incorporating these inequality restrictions.
@@
Moreover, the MDPDE is characterized by a tuning parameter $\gamma$, which controls the trade-off between robustness and efficiency. In the literature, Warwick and Jones 2005 were the first to introduce a data-driven algorithm for determining the optimal value of the tuning parameter, which has since been studied and refined by various authors. Recently, Sugasawa and Yonekura (2021; 2023) delved into this problem by exploiting the Hyvärinen score-matching method with unnormalized models based on robust divergence. Hyvärinen and Dayan 2005 introduced the Hyvärinen score, a score-matching method for continuous-valued data. This approach ensures a consistent estimator by focusing on the gradients of the log of the model and observed data densities eliminating the need to compute the normalization constant. The principle of Hyvärinen score matching (Hyvärinen 2007) involves minimizing the expected squared difference between the score functions of the parameterized density and the true density. However, this approach is restricted to continuous-valued data, where the densities have to be differentiable in the space of $\mathbb{R}^{n}$. For discrete data, Lyu 2012 demonstrated a generalized score-matching approach by incorporating the marginalization operator $\mathcal{M}$. This study was further extended by Xu et al. 2022 for discrete INID multivariate data, where he used a general linear operator $\mathcal{L}$, in place of the gradient operator. In our study, we leverage this generalized score-matching method proposed by Xu et al. 2022 to find the optimal value of the tuning parameter $\gamma.$ Finally, a comparative study is executed to evaluate the performances of an existing method and the generalized score-matching approach.
@@
The rest of the article goes as follows. In Section 2, the model description is provided along with the study of the likelihood function. A review of the minimum density power divergence estimator is presented in Section 3. We derive the restricted minimum density power divergence estimator in Section 4. The determination of optimal tuning parameter $\gamma$ is discussed in Section 5. Section 6 presents an extensive simulation study and real data analysis in order to evaluate the performances of the developed methods.
@@ 2 Model Description
Consider a study involving $m$ independent subjects who may experience $w$ different types of recurrent events. Let $N_{ij}(t)$ denote the number of occurrences of the $j^{th}$ recurrent event up to time $t$ for the $i^{th}$ subject, where $j=1,2,\ldots,w$ and $i=1,\ldots,m$. We assume that for any $i^{th}$ subject, the unobserved heterogeneity is represented by a frailty variable $Z_{i}$, where $Z_{i}^{\prime}s$, for $i=1,\ldots,m$ are independently distributed according to an exponential distribution with a common rate parameter $\zeta(\zeta>0)$. Given $Z_{i}$, $N_{ij}(t)$ are assumed to follow independent nonhomogeneous Poisson process with intensity functions $\lambda_{j}(t\mid Z_{i})=Z_{i}a_{j}t^{b_{j}}$, where $a_{j}>0$, and $b_{j}>0$ for $j=1,\ldots,w$. Further, each $N_{ij}(t)$ is being observed only at finite time points say, $\tau_{1},\tau_{2},\ldots,\tau_{l}$ for $j=1,2,\ldots,w$ and $i=1,\ldots,m$. Therefore, define $N_{ijl}$ as the number of occurrences of the $j^{th}$ recurrent event within the interval $(\tau_{l-1},\tau_{l}]$ for the $i^{th}$ subject, where $l=1,2,\ldots,k$; $j=1,2,\ldots,w$; and $i=1,\ldots,m$. For simplicity, we focus on two recurrent events ($w=2$), although the framework and results can be extended to $w>2$ in a similar manner. Given $Z_{i}$, the probability mass function of $\bm{N_{i}}=(N_{i11},N_{i21},\ldots,N_{i1k},N_{i2k})$ can be derived as follows:
@@
$=\frac{e^{-z_{i}\sum_{j=1}^{2}\sum_{l=1}^{k}{a_{j}(\tau_{l}^{b_{j}}-\tau_{l-1}^{b_{j}})}z_{i}^{\sum_{j=1}^{2}\sum_{l=1}^{k}n_{ijl}}\prod_{j=1}^{2}\prod_{l=1}^{k}(a_{j}(\tau_{l}^{b_{j}}-\tau_{l-1}^{b_{j}}))^{n_{ijl}}}}{\prod_{j=1}^{2}\prod_{l=1}^{k}n_{ijl}!}$ (2.1)
@@
The marginal probability mass function of $\bm{N_{i}}$ can be obtained as,
[…]
@@
[…]
@@
$=\frac{\zeta\prod_{l=1}^{k}(a_{1}(\tau_{l}^{b_{1}}-\tau_{l-1}^{b_{1}}))^{n_{i1l}}(a_{2}(\tau_{l}^{b_{2}}-\tau_{l-1}^{b_{2}}))^{n_{i2l}}\sum_{l=1}^{k}(n_{i1l}+n_{i2l})!}{\prod_{l=1}^{k}n_{i1l}!n_{i2l}!(\zeta+\sum_{l=1}^{k}(a_{1}(\tau_{l}^{b_{1}}-\tau_{l-1}^{b_{1}})+a_{2}(\tau_{l}^{b_{2}}-\tau_{l-1}^{b_{2}})))^{\sum_{l=1}^{k}(n_{i1l}+n_{i2l})}}$ (2.2)
Here, we denote ${\bm{\theta}}=(\zeta,a_{1},b_{1},a_{2},b_{2}).$ Based on the count data across the intervals, the likelihood function can be written as,
@@
$=$ $\frac{\zeta^{m}\prod_{l=1}^{k}(a_{1}(\tau_{l}^{b_{1}}-\tau_{l-1}^{b_{1}}))^{mn_{i1l}}(a_{2}(\tau_{l}^{b_{2}}-\tau_{l-1}^{b_{2}}))^{mn_{i2l}}(\sum_{l=1}^{k}(n_{i1l}+n_{i2l})!)^{m}}{\prod_{l=1}^{k}n_{i1l}^{m}!n_{i2l}^{m}!(\zeta+\sum_{l=1}^{k}(a_{1}(\tau_{l}^{b_{1}}-\tau_{l-1}^{b_{1}})+a_{2}(\tau_{l}^{b_{2}}-\tau_{l-1}^{b_{2}})))^{m\sum_{l=1}^{k}(n_{i1l}+n_{i2l})}}$
Therefore, the log-likelihood can be obtained as,
@@
$-m\sum_{l=1}^{k}(n_{i1l}+n_{i2l})\log(\zeta+\sum_{l=1}^{k}(a_{1}(\tau_{l}^{b_{1}}-\tau_{l-1}^{b_{1}})+a_{2}(\tau_{l}^{b_{2}}-\tau_{l-1}^{b_{2}})))$
Hence, the maximum likelihood estimator (MLE) of ${\bm{\theta}}$, say ${\hat{\bm{\theta}}}$ would be derived as,
${\hat{\bm{\theta}}}=\underset{\bm{\theta}}{\arg\min}\ l({\bm{\theta}})$

### v4

# Inequality Constrained Minimum Density Power Divergence Estimation in Panel Count Data

@@ Abstract
The analysis of panel count data has garnered considerable attention in the literature, leading to the development of multiple statistical techniques. In inferential analysis, most works focus on leveraging estimating equation-based techniques or conventional maximum likelihood estimation. However, the robustness of these methods is largely questionable. In this paper, we present a robust density power divergence estimation method for panel count data arising from non-homogeneous Poisson processes correlated through a latent frailty variable. To cope with real-world incidents, it is often desirable to impose certain inequality constraints on the parameter space, leading to the constrained minimum density power divergence estimator. Being incorporated with inequality restrictions, coupled with the inherent complexity of our objective function, standard computational algorithms are inadequate for estimation purposes. To overcome this, we adopt sequential convex programming, which approximates the original problem through a series of subproblems. Further, we study the asymptotic properties of the resultant estimator, making a significant contribution to this work. The proposed method ensures high efficiency in the model estimation while providing reliable inference despite data contamination. Moreover, the density power divergence measure is governed by a tuning parameter $\gamma$, which controls the trade-off between robustness and efficiency. To effectively determine the optimal value of $\gamma$, this study employs a generalized score-matching technique, marking considerable progress in the data analysis. Simulation studies and real data examples are provided to illustrate the performance of the estimator and to substantiate the theory developed.
@@
Asymptotic property , Density power divergence , Inequality constraints , Optimal tuning parameter , Robustness
journal: Applied Mathematical Modelling
@@
{highlights}
Proposes a robust estimator for panel count data with latent frailty.
Incorporates inequality constraints into density power divergence estimation.
Establishes asymptotic properties under parameter restrictions.
Finds optimal tuning parameter through generalized score-matching technique.
Validates performance through simulations and real applications.
@@ 1 Introduction
In the study of event history, the types of events that occur repetitively are usually referred to as recurrent events. Some examples of recurrent events refer to hospitalizations of intravenous drug users, multiple instances of specific tumors [1], repeated pyogenic infections in patients with inherited disorders [2], and claims for vehicle warranties [3]. Panel count data are common when subjects in a study encounter recurrent events observed only at discrete time points instead of being observed continuously in time. They might occur when continuous monitoring of the subjects under study is too expensive or infeasible. As a result, panel count data only provide the number of occurrences of the recurrent events of interest between successive observation times, while the exact occurrence times remain unknown; refer to Huang et al. [4], Sun and Zhao [5].
@@
In the literature, it is a common practice to characterize the occurrences of recurrent events by leveraging counting processes. Further, due to the incomplete nature of the panel count data, it is more convenient to model the mean function of the counting processes. Readers may refer to Sun and Kalbfleisch [6], Wellner and Zhang [7], Hu et al. [8], and references therein. In this work, we focus on the analysis of panel count data where each subject under study is exposed to multiple recurrent events characterized by nonhomogeneous Poisson processes; refer to Tang et al. [9], Li and Pham [10]. It is a non-stationary counting process with independent increments. In many instances, as it proves challenging to assess all the covariates, a latent frailty variable can be considered that reflects individual-specific heterogeneity typically influencing the frequency of these recurrent incidents; follow Slimacek and Lindqvist [11]. In this work, we develop the inferential analysis of panel count data where each subject under study is exposed to multiple recurrent events characterized by nonhomogeneous Poisson processes correlated through a frailty variable. Based on the counting process, there exists a substantial body of literature concerning inferential studies for panel count data. Amongst them, He et al. [12] exploited a semi-parametric approach for the analysis of panel count data. Moreover, Sun and Wei [13] and He et al. [14] investigated regression analysis based on counting processes concerning univariate and multivariate panel count data, respectively.
@@
Throughout the years, most research on estimation methods for panel count data has been devoted to maximum likelihood estimation (MLE). The major contributions consist of the works by Wellner and Zhang [15], Zhu et al. [16], Li et al. [17], Zeng and Lin [18]. Maximum likelihood estimation is a popular method because of its recognized characteristics, such as asymptotic efficiency, consistency, sufficiency, and invariance. However, with small deviations from the assumed model conditions, those methods often produce biased and inefficient estimates, highlighting the necessity for robust estimation techniques. In this paper, our objective is to perform statistical inference based on multivariate panel count data by applying robust density power divergence estimation methods. The minimum density power divergence estimation (MDPDE) was first proposed by Basu et al. [19] and subsequently applied in numerous studies. Readers can track the publications of Ghosh and Basu [20], Basu et al. [21], Mandal et al. [22], Baghel and Mondal [23]. Recently, Balakrishnan et al. [24] developed weighted minimum density power divergence estimators to analyze one-shot device testing data, similar to current status data, which is a specific type of panel count data. Xiong and Zhu [25] developed a robust MDPDE-based method for negative binomial integer-valued GARCH models, effectively handling overdispersed count time series data and outliers. Their research motivates our study on panel count data, where robustness and heterogeneity are key concerns. Upon reviewing the literature, it is found that none of the existing studies on multivariate panel count data generated from the counting process address robust methods for parameter estimation. Hence, we suggest this novel approach to serve the need for a robust estimation technique.
@@
Further, in real-life scenarios, it is often observed that one recurring event takes place more frequently than the other event. Such instances can be accurately characterized by imposing restrictions on the parameter space. Moreover, research on restricted minimum density power divergence estimators has been sparse, with all current studies concentrating solely on equality constraints. For example, Basu et al. [26], Felipe et al. [27], Baghel and Mondal [28] suggested equality constraints in the parameter space of the model parameters and derived the restricted minimum density power divergence estimators. In this study, we broaden the scope of restricted minimum density power divergence estimators by introducing inequality constraints into the parameter space. However, the addition of these constraints, along with the complex nature of the objective function, limits the effectiveness of standard computational algorithms in finding optimal solutions. This challenge has motivated our exploration of Sequential Convex Programming (SCP), a technique initially developed by Dinh and Diehl [29], which transforms a non-convex problem into a sequence of convex subproblems. A key component of SCP is the trust region, which constrains each subproblem to a neighborhood around the current iterate, ensuring both convergence and local validity of convex approximations. SCP has found successful applications in trajectory optimization, model predictive control, and various engineering design problems. For additional insights and applications, readers are encouraged to consult Debrouwere et al. [30], Morgan et al. [31], and Wang and Lu [32], among others. The principal novelty in this paper lies in deriving the asymptotic distribution of the minimum density power divergence estimators while incorporating these inequality restrictions.
@@
Moreover, the MDPDE is characterized by a tuning parameter $\gamma$, which controls the trade-off between robustness and efficiency. In the literature, Warwick and Jones [33] were the first to introduce a data-driven algorithm for determining the optimal value of the tuning parameter, which has since been studied and refined by various authors. Recently, Sugasawa and Yonekura [34], [35] delved into this problem by exploiting the Hyvärinen score-matching method with unnormalized models based on robust divergence. Hyvärinen and Dayan [36] introduced the Hyvärinen score, a score-matching method for continuous-valued data. This approach ensures a consistent estimator by focusing on the gradients of the log of the model and observed data densities, eliminating the need to compute the normalization constant. The principle of Hyvärinen score matching (Hyvärinen [37]) involves minimizing the expected squared difference between the score functions of the parameterized density and the true density. However, this approach is restricted to continuous-valued data, where the densities have to be differentiable in the space of $\mathbb{R}^{n}$. For discrete data, Lyu [38] demonstrated a generalized score-matching approach by incorporating the marginalization operator $\mathcal{M}$. This study was further extended by Xu et al. [39] for discrete independent multivariate data, where he used a general linear operator $\mathcal{L}$, in place of the gradient operator. In our study, we leverage this generalized score-matching method proposed by Xu et al. [39] to find the optimal value of the tuning parameter $\gamma.$ Finally, a comparative study is executed to evaluate the performance of an existing method and the generalized score-matching approach.
@@
The rest of the article goes as follows. In Section 2, the model description is provided along with the study of the likelihood function. A review of the minimum density power divergence estimator is presented in Section 3. We derive the restricted minimum density power divergence estimator in Section 4. The determination of optimal tuning parameter $\gamma$ is discussed in Section 5. Section 6 presents an extensive simulation study and real data analysis in order to evaluate the performance of the developed methods.
@@ 2 Model Description
Consider a study involving $m$ independent subjects who may experience $w$ different types of recurrent events. Let $N_{ij}(t)$ denote the number of occurrences of the $j^{th}$ recurrent event up to time $t$ for the $i^{th}$ subject, where $j=1,2,\ldots,w$ and $i=1,\ldots,m$. We assume that for any $i^{th}$ subject, the unobserved heterogeneity is represented by a frailty variable $Z_{i}$, where $Z_{i}^{\prime}s$, for $i=1,\ldots,m$ are independently distributed according to an exponential distribution with a common rate parameter $\zeta(\zeta>0)$. Given $Z_{i}$, $N_{ij}(t)$ are assumed to follow independent nonhomogeneous Poisson process with intensity functions $\lambda_{j}(t\mid Z_{i})=Z_{i}a_{j}t^{b_{j}}$, where $a_{j}>0$, and $b_{j}>0$ for $j=1,\ldots,w$. Further, each $N_{ij}(t)$ is being observed only at finite time points say, $\tau_{1},\tau_{2},\ldots,\tau_{l}$ for $j=1,2,\ldots,w$ and $i=1,\ldots,m$. Therefore, define $N_{ijl}$ as the number of occurrences of the $j^{th}$ recurrent event within the interval $(\tau_{l-1},\tau_{l}]$ for the $i^{th}$ subject, where $l=1,2,\ldots,k$; $j=1,2,\ldots,w$; and $i=1,\ldots,m$. For simplicity, we focus on two recurrent events ($w=2$), although the framework and results can be extended to $w>2$ in a similar manner. Given $Z_{i}$, the probability mass function of $\bm{N_{i}}=(N_{i11},N_{i21},\ldots,N_{i1k},N_{i2k}$ can be derived as follows
@@
$=\frac{e^{-z_{i}\sum_{j=1}^{2}\sum_{l=1}^{k}{a_{j}(\tau_{l}^{b_{j}}-\tau_{l-1}^{b_{j}})}z_{i}^{\sum_{j=1}^{2}\sum_{l=1}^{k}n_{ijl}}\prod_{j=1}^{2}\prod_{l=1}^{k}(a_{j}(\tau_{l}^{b_{j}}-\tau_{l-1}^{b_{j}}))^{n_{ijl}}}}{\prod_{j=1}^{2}\prod_{l=1}^{k}n_{ijl}!},$ (1)
@@
The marginal probability mass function of $\bm{N_{i}}$ can be obtained as
$f(\bm{n_{i}})$ $=\int_{0}^{\infty}P(\bm{N_{i}}=\bm{n_{i}}|\ z_{i})h(z_{i})~dz_{i}$
@@
$=\int_{0}^{\infty}\frac{e^{-z_{i}\sum_{j=1}^{2}\sum_{l=1}^{k}{a_{j}(\tau_{l}^{b_{j}}-\tau_{l-1}^{b_{j}})}z_{i}^{\sum_{j=1}^{2}\sum_{l=1}^{k}n_{ijl}}\prod_{j=1}^{2}\prod_{l=1}^{k}(a_{j}(\tau_{l}^{b_{j}}-\tau_{l-1}^{b_{j}}))^{n_{ijl}}}}{\prod_{j=1}^{2}\prod_{l=1}^{k}n_{ijl}!}\zeta e^{-\zeta z_{i}}~dz_{i}$
@@
$=\frac{\zeta\prod_{l=1}^{k}(a_{1}(\tau_{l}^{b_{1}}-\tau_{l-1}^{b_{1}}))^{n_{i1l}}(a_{2}(\tau_{l}^{b_{2}}-\tau_{l-1}^{b_{2}}))^{n_{i2l}}\sum_{l=1}^{k}(n_{i1l}+n_{i2l})!}{\prod_{l=1}^{k}n_{i1l}!n_{i2l}!(\zeta+\sum_{l=1}^{k}(a_{1}(\tau_{l}^{b_{1}}-\tau_{l-1}^{b_{1}})+a_{2}(\tau_{l}^{b_{2}}-\tau_{l-1}^{b_{2}})))^{\sum_{l=1}^{k}(n_{i1l}+n_{i2l})}}.$ (2)
Here, we denote ${\bm{\theta}}=(\zeta,a_{1},b_{1},a_{2},b_{2}).$ Based on the count data across the intervals, the likelihood function can be written as
@@
$=$ $\frac{\zeta^{m}\prod_{l=1}^{k}(a_{1}(\tau_{l}^{b_{1}}-\tau_{l-1}^{b_{1}}))^{mn_{i1l}}(a_{2}(\tau_{l}^{b_{2}}-\tau_{l-1}^{b_{2}}))^{mn_{i2l}}(\sum_{l=1}^{k}(n_{i1l}+n_{i2l})!)^{m}}{\prod_{l=1}^{k}n_{i1l}^{m}!n_{i2l}^{m}!(\zeta+\sum_{l=1}^{k}(a_{1}(\tau_{l}^{b_{1}}-\tau_{l-1}^{b_{1}})+a_{2}(\tau_{l}^{b_{2}}-\tau_{l-1}^{b_{2}})))^{m\sum_{l=1}^{k}(n_{i1l}+n_{i2l})}}.$
Therefore, the log-likelihood can be obtained as
@@
$-m\sum_{l=1}^{k}(n_{i1l}+n_{i2l})\log(\zeta+\sum_{l=1}^{k}(a_{1}(\tau_{l}^{b_{1}}-\tau_{l-1}^{b_{1}})+a_{2}(\tau_{l}^{b_{2}}-\tau_{l-1}^{b_{2}}))).$
Hence, the maximum likelihood estimator (MLE) of $\bm{\theta}$}, say $^{\hat{\bm{\theta}}}$ would be derived as
${\hat{\bm{\theta}}}=\underset{\bm{\theta}}{\arg\min}\ l({\bm{\theta}}).$